OpenStudy (anonymous):

Determine wheather he grapghs of he given of the given equations are parallel perpendicular, or neither. explian. 1. y = 4x + 5 _4x + y = -13 2. 3x + 6y = 12 y - 4 = -1/2 (fraction) (x + 2)

OpenStudy (mertsj):

Can you find the slopes of these lines?

OpenStudy (mertsj):

It is very difficult to help you if you won't respond.

OpenStudy (anonymous):

i dnt no how dat much

OpenStudy (mertsj):

y=4x+5 has slope 4 -4x+y = -13 is the same as y = 4x-13. It too has slope 4 Since the two lines have the same slope, they are parallel.

OpenStudy (anonymous):

ok so the anwser is y+ax-13

OpenStudy (mertsj):

3x+6y=12 is the same as y=-1/2x+2 y-4=-1/2(x-2) is the same as y = -1/2x+5 The slope of both of these lines is -1/2 so they too are parallel.

OpenStudy (mertsj):

What does the problem ask you to do?

OpenStudy (anonymous):

it ask me to Determine wheather he grapghs of he given of the given equations are parallel perpendicular, or neither. explian.

OpenStudy (mertsj):

Well then, are they parallel or perpendicular or neither?

OpenStudy (anonymous):

r u talking about the first one or the secoung one

OpenStudy (mertsj):

The first one.

OpenStudy (anonymous):

parelle

OpenStudy (mertsj):

How do you know?

OpenStudy (anonymous):

i dont thts y i asked the questtion

OpenStudy (mertsj):

I don't know what that means.

OpenStudy (mertsj):

The answer to number 1 is parallel because they have the same slope. The answer to number 2 is parallel because they have the same slope.